OpenStudy (kyanthedoodle):

Explain the law of motion needed to solve this problem and then solve it. On Planet Zorg, a 30 kg barbell can be lifted by only exerting a force of 180 N. What is the acceleration of gravity on Planet Zorg? Predict how the motion of this object would change if on Earth. Please help. I have no idea how to do this.

OpenStudy (elonasushchik):

From Newton's second law (F=ma): a = F/m = 189 N/30 kg = 6.0 m/s2

OpenStudy (kyanthedoodle):

Thanks man.